OpenStudy (anonymous):

4(1/2x+7)-3x=1/3(9-6)-23

OpenStudy (anonymous):

\[4\left( \frac{ 1}{ 2x + 7 }\right)-3x=\frac{ 1 }{ 3(9-6)-23 }\]

OpenStudy (anonymous):

is that your question?..

OpenStudy (anonymous):

no its 4(.5x+7)-3x=.33(9-6)-23 that .5 is 1over 2 (1/2 and that .3 is 1over 3 (1/3)

OpenStudy (ajprincess):

 Can u do nw?